the area of a square is 32 square units. if a side length is x+2.5 find x to the nearest tenth
Answer by Boreal(15235)

Area is 32 sq units. It is a square, so sides are equal and area is s^2
s^2=32
s=4 sqrt (2) ;; sqrt(16)* sqrt (2)
That is 5.66 or 5.7 to the nearest tenth.
BUT, the side is equal to X+2.5
Therefore, X+2.5=5.7
X=3.2
Can do it by setting x^2+5x +6.25=32; (x+2.5)^2 expanded
x^2+5x-25.75=0 and using the quadratic formula
Positive square root only
[-5 + sqrt (25+103}/2
=[-5 +sqrt (128)]/2
(1/2) (-5+11.3)= (1/2)(6.3)=3.2 with rounding